# On a Super-analogue of Carrollian Manifolds

### v2

@@ Abstract.
Inspired by Carrollian geometry, we define super-Carrollian manifolds as a supermanifold with an even degenerate metric such that the kernel is generated by a non-singular odd vector field that is a supersymmetry generator. Alongside other results, we show that compatible affine connections always exist, albeit they must carry torsion. As a physically relevant example, we show how a super-Carrollian manifold can be constructed from standard superspace $\mathbb{R}^{4|4}$ via an Inönü–Wigner contraction.
@@ 1. Introduction
Carrollian manifolds are understood as manifolds equipped with a degenerate metric whose kernel is spanned by a nowhere vanishing complete vector field (see [10, 11, 12, 19, 28, 29]). Carrollian geometry/physics has grown from a mathematical curiosity based on the ultra-relativistic limit $c\rightarrow 0$ to a subject of ongoing research. For a review of many of the facets of Carrollian physics, the reader may consult Bagchi et al. [2]. Intrinsic approaches to Carrollian geometry, so working with geometries not directly associated with the ultra-relativistic limit, have been developed, starting with the work of Duval et al. [10, 11, 12]. Null hypersurfaces, such as punctured future or past light-cones in Minkowski spacetime, and the event horizon of a Schwarzschild black hole, are examples of Carrollian manifolds.
@@
Supermanifolds offer the possibility of generalisations of classical geometries that are described by odd structures. For example, the notion of odd connections on supermanifolds was studied by the author & Grabowski [8]; Khudaverdian & Peddie [22] provided a comparison between odd Riemannian and odd symplectic; and Khudaverdian & Voronov [23] examined odd Laplace operators. Thus, as a mathematical question, the possibility of Grassmann odd analogues of Carrollian geometries is raised. Specifically, one can consider an odd vector field that generates the kernel of a degenerate metric. Shander [30] provides the local form of the (non-singular) odd vector fields
Homological: $Q=\frac{\partial}{\partial\tau}$, $Q^{2}=\mathbf{0}$, Supersymmetric: $Q=\frac{\partial}{\partial\tau}+\tau\frac{\partial}{\partial t}$, $Q^{2}\neq\mathbf{0}$,
@@
where $t$ is an even coordinate and $\tau$ an odd coordinate. It is the supersymmetric generalisation that we study here, as the non-integrable distribution is quite at odds with the classical situation. Moreover, Vaintrob [31] has shown that if a supermanifold admits a non-singular homological vector field, then the supermanifold is a trivial odd line bundle. We will restrict our attention to even degenerate metrics, although odd metrics can be studied; they seem less relevant in physics [^1] The author is not aware of any direct application of odd Riemannian metrics in physics. This is in contrast to the use of odd symplectic structures in the BV formalism. See Khudaverdian [21] and references therein..
@@
The core concept introduced in the paper is that of a super-Carrollian manifold. That is, a supermanifold with a single odd direction that is equipped with a degenerate (even) Riemannian metric such that the kernel of the metric is generated by a non-singular odd vector field $Q$, that satisfies $Q^{2}\neq\mathbf{0}$; see Definition 2.12. In particular, $[Q,Q]=2P$, and $[P,Q]=0$ and so we have the $N=1$, $d=1$ supertranslation algebra describing the kernel of the degenerate Riemannian metric.
@@
We define the super-Carrollian Lie algebra as the Lie superalgebra of infinitesimal automorphisms of a super-Carrollian manifold. One observation is that this Lie superalgebra is pure even if $P$ is not Killing, and if $P$ is Killing, the odd elements are spanned by the supersymmetric covariant derivative. Affine connections on a super-Carrollian manifold are carefully studied. Supersymmetric and metric compatible affine connections are defined and examined; the first are connections for which the vector field $Q$ is parallel, and the latter is the natural generalisation of metric compatible connections to degenerate metrics. As the metric on a super-Carrollian manifold is degenerate, there is no direct generalisation of the fundamental theorem of Riemannian supergeometry. We prove that both supersymmetric compatible and metric compatible connections exist on any super-Carrollian manifold; however, they are not uniquely fixed by the metric and odd vector field. Moreover, due to the distribution given by the kernel of the metric being non-integrable, these connections must carry torsion.
@@
Starting with flat superspace one can construct a super-Carrollian manifold using an ultra-relativistic limit. Detailed definitions and presentation of conventions can be found throughout Subsections 2.1 and 2.2. We will follow the conventions of [7] for spinors and gamma matrices. Consider $N=1$, $d=4$ flat superspace ${\mathbb{R}}^{4|4}$, with global coordinates $(x^{\mu},\theta_{\alpha})$, where the odd coordinates are real Majorana spinors. The supersymmetry generators are
@@
To take a Carrollian limit (Inönü–Wigner contraction), we separate spatial and temporal directions and write
$(x^{\mu},\theta_{\alpha})=(x^{i},t;\,\theta_{a},\tau),$
@@
where $c$ is the speed of light, regarded as a contraction parameter. The temporal coordinates $(t,\tau)$ are not rescaled. Furthermore, we resale the generators
@@
to ensure that the (graded) commutators remain finite in the ultra-relativistic limit. In the limit $c\to 0$, the rescaled generators vanish except for the surviving pair
@@
$g=\textnormal{d}x^{i}\otimes\textnormal{d}x^{j}\,\delta_{ji}+2\,\textnormal{d}t\otimes\textnormal{d}\tau\,\tau-\textnormal{d}t\otimes\textnormal{d}t\,.$
@@
$\langle Q|\partial_{i}\rangle=\langle\partial_{\tau}|\partial_{i}\rangle+\tau\,\langle\partial_{t}|\partial_{i}\rangle=0\,$,
@@
which demonstrates that the kernel of $g$ is generated by $Q$. Notice that the reduce metric, so informally setting $\tau=0$, gives the Minkowski metric on ${\mathbb{R}}^{4}$.
The above constructions will from the basis of examples 2.30 and 2.43.
The general definition of a super-Carrollian manifold (see Definition 2.12) covers the above example, however, there will be no assumption that the supermanifold necessarily arises from a Inönü–Wigner contraction. That is, we take an intrinsic point of view.
@@
  • Main Theorem - affine connections that are both supersymmetric compatible and metric compatible exist on any super-Carrollian manifold;
@@
  • Proposition 2.18 - the reduced metric is non-degenerate (provided the number of even coordinates is not $2$);
@@
  • Proposition 2.25 + Corollary 2.26 - the super-Carrollian Lie algebra is finite dimensional (provided the number of even coordinates is not 22).
The case of two even coordinates requires careful attention. In particular, the reduced metric, so the induced structure on the underlying smooth manifold, is not guaranteed to be non-degenerate. The question of degeneracy needs to be addressed case by case. The reader may consult examples 2.16 and 2.17 for a demonstration of this fact. Proving that the super-Carrollian Lie algebra is finite critically depends on the reduced metric being non-degenerate.
Arrangement. In Section 2 we proceed with the bulk of this paper. We recall the fundamental theory of supermanifolds equipped with (degenerate) metrics and connections in Subsection 2.1. In Subsection 2.2, the concept of a super-Carrollian manifold is presented, and some immediate results are given. The super-Carrollian Lie algebra is studied in Subsection 2.3. The question of compatible affine connections is addressed in Subsection 2.4. We end in Section 3 with some concluding remarks.
@@ 2.1. Even Metrics and Connections on Supermanifolds
We will assume that the reader has some familiarity with the category of (real and finite-dimensional) supermanifolds SMan. We understand a supermanifold $M:=(|M|,\>\mathcal{O}_{M})$ of dimension $n|m$ to be a supermanifold as defined by Berezin & Leites [5, 27], i.e., we take the locally ringed space approach. For an overview of the general theory of supermanifolds, the reader may consult, for example, [9, 26, 32]. Underlying any supermanifold is a smooth manifold that we will denote $M_{red}=(|M|,C^{\infty}_{|M|}(-))$. An incomplete list of works on Riemannian supermanifolds includes [13, 14, 15, 16, 17, 24]. The warning from the outset is that we will include degenerate metrics in our definition.
@@
for all (homogeneous) $X,Y,Z\in{Vect}(M)$ and $f\in C^{\infty}(M)$. Extension of these properties to inhomogeneous vector fields is by linearity.
@@
A metric $g$ is said to be non-degenerate if $\ker(g)=\{\mathbf{0}\}$, and is said to be degenerate otherwise.
@@
Killing vector fields are defined in exactly the same way as in classical Riemannian geometry.
@@
This follows in complete parallel with the classical case using $L_{[X,Y]}=[L_{X},L_{Y}]$. ∎
@@
Affine connections exist on any (real) supermanifold. This can be proved by adapting the standard arguments, i.e., affine connections are local operators and the existence of a partition of unity [^2] See [27, Lemma 3.1.7 and Corollary 3.1.8] for the existence of partitions of unity and bump functions on supermanifolds..
@@
2.2. Super-Carrollian Manifolds
@@
A super-Carrollian manifold is a quadruple $(M,g,Q,P)$, where
@@
subject to the compatibility condition
$\ker(g)=\operatorname{Span}\{Q\}\,.$
A morphism of super-Carrollian manifolds $\Phi:(M,g,Q,P)\rightarrow(M^{\prime},g^{\prime},Q^{\prime},P^{\prime})$ is a diffeomorphism $\Phi:M\rightarrow M^{\prime}$, such that
@@
The resulting category of super-Carrollian manifolds is denoted SCarMan. The group of automorphism of a super-Carrollian manifold we denote as $\mathrm{SCarr}(M,g,Q,P)$, and the associated super-Carrollian Lie algebra $\mathfrak{scarr}(M,g,Q,P)$ is the Lie superalgebra of vector fields $X\in{Vect}(M)$ that satisfy
@@
    the rank of the kernel is $0|1$, and the constant rank ensures that the kernel is well-defined as a locally free module;
@@
    the supermanifold $M$ cannot have more than one ‘odd direction’ as they are automatically null directions and would violate the rank condition. Moreover, $M$ is an odd line bundle, i.e., locally $M$ is of the form $|U|\times{\mathbb{R}}^{0|1}$;
@@
    Shander’s theorem [30] means that in the neighbourhood of any point $m\in|M|$, adapted coordinate can always be found, which we will refer to as Shander coordinates, $(x^{a},t,\tau)$ such the vector fields take the canonical form
@@
    morphisms equate the kernels, i.e., $\Phi_{*}\ker(g)=\ker(g^{\prime})$, and $\Phi_{*}P=P^{\prime}$ as the push-forward by a diffeomorphism is a homomorphism of Lie brackets;
@@
    if $X\in\mathfrak{scarr}(M,g,Q,P)$, then $\mathcal{L}_{X}P=[[X,Q],Q]=0$.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold. Then in Shander coordinates $(x^{a},t,\tau)$, the most general from of the degenerate metric $g$ is
@@
In Shander coordinates $(x^{a},t,\tau)$ the most general form of a (possibly degenerate) metric on $M$ is of the form
@@
Locally, $\tau Q=\tau\partial_{\tau}$, and so these vector fields are not linearly independent. Thus, $\langle Q|\partial_{\tau}\rangle=0$ is consistent with the rank of $\ker(g)$ being generated by $Q$.
@@
which apriori, may be a degenerate. Using coordinates on $M_{red}$ induced by Shander coordinates $(x^{a},t,\tau)$, the metric $g_{red}$ is of the form
@@
Example 2.15.
Consider $M={\mathbb{R}}^{1|1}$ equipped with global coordinates $(t,\tau)$. Degenerate metrics on ${\mathbb{R}}^{1|1}$ of the form
where $g_{tt}$ is a smooth function $t$. To have a constant rank kernel, it must be the case that $g_{tt}$ is nowhere vanishing on $\mathbb{R}$}. The reader can quickly check that we do indeed have a super-Carrollian manifold, i.e., the kernel of $g$ is generated by $Q$. The reduced metric is $g_{red}=\textnormal{d}t\otimes\textnormal{d}t\,g_{tt}(t)$, which is non-degenerate.
@@
The reader can quickly check that we do indeed have a super-Carrollian manifold, i.e., the kernel of $g$ is generated by $Q$. The reduced metric is $g_{red}=\textnormal{d}x\otimes\textnormal{d}x+\textnormal{d}t\otimes\textnormal{d}t$, which is non-degenerate.
Example 2.17.
$g=\textnormal{d}x\otimes\textnormal{d}x+2\,\textnormal{d}x\otimes\textnormal{d}t-2\;\textnormal{d}x\otimes\textnormal{d}\tau\,\tau-2\,\textnormal{d}t\otimes\textnormal{d}\tau\,\tau+\textnormal{d}t\otimes\textnormal{d}t\,.$
The reader can quickly check that we do indeed have a super-Carrollian manifold, i.e., the kernel of $g$ is generated by $Q$. The reduced metric is $g_{red}=\textnormal{d}x\otimes\textnormal{d}x+2\,\textnormal{d}x\otimes\textnormal{d}t+\textnormal{d}t\otimes\textnormal{d}t$. To check the degeneracy, written as a matrix
$g_{red}=\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ $\det(g_{red})=0$ $\Longrightarrow~~\textnormal{non-invertable}$
Thus, $g_{red}$ is degenerate.
The examples above suggest that the reduced metric, in general, may be degenerate. However, as we will argue, the potential degeneracy in $g_{red}$ is restricted to the $n=2$ case.
@@
Let $(M,g,Q,P)$ be a super-Carrollian manifold such that $\dim M_{red}\neq 2$. Then the reduced manifold $M_{red}$ is a pseudo-Riemannian manifold. If $\dim M_{red}=2$, then reduced metric on $M_{red}$ maybe degenerate.
@@
As the question of degeneracy can be addressed locally, we will employ Shander coordinates $(x^{a},t,\tau)$, so that $Q=\partial_{\tau}+\tau\partial_{t}$ and $P=\partial_{t}$. By Proposition 2.13 the reduced metric has the form
@@
As a block matrix we have
$g_{red}=\begin{pmatrix}g_{ab}&g_{at}\\[4.0pt] g_{tb}&g_{tt}\end{pmatrix}\,$,
where the entries are ordinary smooth functions in $x^{a}$ and $t$.
Whenever $g_{ab}$ is invertible we may factor the determinant by the Schur complement:
$\det(g_{red})=\det(g_{ab})\cdot S,\qquad S:=g_{tt}-g_{ta}\,(g_{ab})^{-1}g_{bt}\,.$
@@
Claim 1: for $n>2$, $\det(g_{ab})\neq 0$.
We will prove this via a contradiction. Let us assume that there exists a $U=U^{a}\partial_{a}$ such that $U^{a}g_{ab}=0$. We remark that if $n=2$, then $g_{ab}$ is a single function and this may be zero while not violating the condition on the rank of $g$. Then any vector field $X=X^{a}\partial_{a}+X^{t}\partial_{t}+X^{\tau}\partial_{\tau}$ we have
@@
Next consider a new vector field $U^{\prime}=U+\lambda\,P$, where $\lambda$ is an even function. Then
@@
Claim 2: for $n>2$, The Schur scalar $S$ is nonzero.
Claim 1 establishes that $g_{ab}$ is invertible and we will denote the inverse as $g^{ab}$, as standard. The Schur scalar is $S=g_{tt}-g_{ta}g^{ab}g_{bt}$. If $S=0$, then there exists a vector field $V\in\ker(g_{red})$. Every such even vector field has a canonical lift, also denoted here by $V$, as it does not depend on the odd coordinate. Thus, $\epsilon(\langle V|X\rangle)=0$ for all lifts $X$, and so $\langle V|X\rangle=0$. Thus, $V$ is in the kernel of $g$. However, this is in violation of the dimensions of the rank of the kernel. This implies it must the the case that $S\neq 0$.
Thus, for $n>2$, $\det(g_{red})\neq 0$ and so $g_{red}$ defines a pseudo-Riemannian structure on $M_{red}$.
The case of $n=1$ is covered by Example 2.15. The reduced metric $g_{red}$ is always non-degenerate. ∎
Corollary 2.19.
Via Proposition 2.13, the degenerate metric $g$ of a super-Carrollian manifold is determined $g_{red}$.
@@
where $g_{\mathbb{R}}$ is the constant metric on $\mathbb{R}$}. Then $M=M_{0}\times{\mathbb{R}}^{1|1}$ is a super-Carrollian manifold with the degenerqte metric in Shander coordinates being of the form
@@
2.3. The Super-Carrollian Lie Algebra
The super-Carrollian Lie algebra was defined earlier as the infinitesimal automorphisms of a super-Carrollian manifold (see Definition 2.12). The Lie superalgebra $\mathfrak{scarr}(M,g,Q,P)$ is defined as the Lie superalgebra of vector fields $X\in{Vect}(M)$ that satisfy $\mathcal{L}_{X}g=0$ and $\mathcal{L}_{X}Q=0$. In this subsection, we examine the structure of this Lie superalgebra and establish that it is finite dimensional.

### v3

# On a Grassmann odd analogue of Carrollian Manifolds

@@ Abstract.
We define a Grassmann odd analogue of a Carrollian manifold as a supermanifold of dimension $n|1$ with an even degenerate metric such that the kernel is generated by a non-singular odd vector field that is a supersymmetry generator. Alongside other results, we establish that the reduced manifold is a pseudo-Riemannian manifold, and show that compatible affine connections always exist, albeit they must carry torsion. As a physically relevant example, we examine an Inönü–Wigner contraction of the supertranslation algebra on standard superspace $\mathbb{R}^{4|4}$.
@@ 1. Introduction
Carrollian manifolds are understood as manifolds equipped with a degenerate metric whose kernel is spanned by a nowhere vanishing complete vector field (see Duval et al. [9, 10, 11], and the earlier works of Lévy-Leblond [27], Sen Gupta [29], and Henneaux [19]). Carrollian geometry/physics has grown from a mathematical curiosity based on the non-relativistic limit $c\rightarrow 0$ to a subject of ongoing research. For a review of many of the facets of Carrollian physics, the reader may consult Bagchi et al. [1]. Intrinsic approaches to Carrollian geometry, so working with geometries not directly associated with a limit, have been developed, starting with the work of Duval et al. [9, 10, 11] (also see Henneaux [19]). Null hypersurfaces, such as punctured future or past light-cones in Minkowski spacetime, and the event horizon of a Schwarzschild black hole, are examples of Carrollian manifolds.
@@
Supermanifolds offer the possibility of generalisations of classical geometries that are described by odd structures. For example, the notion of odd connections on supermanifolds was studied by the author & Grabowski [7]; Khudaverdian & Peddie [22] provided a comparison between odd Riemannian and odd symplectic; and Khudaverdian & Voronov [23] examined odd Laplace operators. Thus, as a mathematical question, the possibility of Grassmann odd analogues of Carrollian geometries is raised. Specifically, one can consider an odd vector field that generates the kernel of a degenerate metric. Shander [30] provides the local form of the (non-singular) odd vector fields Homological: $Q=\frac{\partial}{\partial\tau}$, $Q^{2}=\mathbf{0}$, Supersymmetric: $Q=\frac{\partial}{\partial\tau}+\tau\frac{\partial}{\partial t}$, $Q^{2}=\frac{\partial}{\partial t}$,
@@
where $t$ is an even coordinate and $\tau$ an odd coordinate. It is the supersymmetric generalisation that we study here, as the non-integrable distribution is quite at odds with the classical situation. Moreover, Vaintrob [31] has shown that if a supermanifold admits a non-singular homological vector field, then the supermanifold is a trivial odd line bundle. We will restrict our attention to even degenerate metrics, although odd metrics can be studied; they seem less relevant in physics [^1] The author is not aware of any direct application of odd Riemannian metrics in physics; this explains the general lack of their appearance in the literature. This is in contrast to the use of odd symplectic structures in the BV formalism. See Khudaverdian [21] and references therein..
@@
The core concept introduced in the paper is that of a super-null Riemannian manifold. That is, a supermanifold with a single odd direction that is equipped with a degenerate (even) Riemannian metric $g$, such that the kernel of the metric is generated by a non-singular odd vector field $Q$, that satisfies $Q^{2}\neq\mathbf{0}$; see Definition 2.9. In particular, $[Q,Q]=2P$, and $[P,Q]=0$, and so we have the $N=1$, $d=1$ supertranslation algebra describing the kernel of the degenerate Riemannian metric. We avoid the nomenclature “super-Carrollian manifolds” or similar to avoid confusion with supersymmetric extensions of Carrollian geometry and/or the Carroll group. We stress that the structures we study are not associated with Carrollian superalgebras as studied in the context of Carrollian supergravities, see for example, Bergshoeff, Gomis, & Parra [4], Ravera & Zorbra [28], and Grumiller, Montecchio, & Nejati [16], and references therein.
@@
We define the super-null Lie algebra as the Lie superalgebra of infinitesimal automorphisms of a super-null Riemannian manifold, i.e., vector fields $X$ that satisfy $\mathcal{L}_{X}g=0$ and $\mathcal{L}_{X}Q=0$. One observation is that this Lie superalgebra is pure even if $P$ is not Killing, and if $P$ is Killing, the odd elements are spanned by the supersymmetric covariant derivative. Affine connections on a super-null Riemannian manifold are carefully studied. Supersymmetric and metric compatible affine connections are defined and examined; the former are connections for which the vector field $Q$ is parallel, and the latter are the natural generalisation of metric compatible connections to degenerate metrics. As the metric on a super-null Riemannian manifold is degenerate, there is no direct generalisation of the fundamental theorem of Riemannian supergeometry. We prove that both supersymmetric compatible and metric compatible connections exist on any super-null Riemannian manifold; however, they are not uniquely fixed by the metric and odd vector field. Moreover, due to the distribution given by the kernel of the metric being non-integrable, these connections must carry torsion.
@@
Starting with flat superspace, one can construct a super-null Riemannian manifold using an Inönü–Wigner contraction (see [20]). Detailed definitions and presentation of conventions can be found throughout Subsections 2.1 and 2.2. We will follow the conventions of [6] for spinors and gamma matrices. Consider $N=1$, $d=4$ flat superspace ${\mathbb{R}}^{4|4}$, with global coordinates $(x^{\mu},\theta_{\alpha})$, where the odd coordinates are real Majorana spinors. The supersymmetry generators are
@@
To take an Inönü–Wigner contraction, we separate spatial and temporal directions and write
$(x^{\mu},\theta_{\alpha})=(x^{i},t,\theta_{a},\tau)\,.$
@@
where $c$ is the speed of light, regarded as a contraction parameter. The temporal coordinates $(t,\tau)$ are not rescaled. Furthermore, we rescale the generators
@@
to ensure that the (graded) commutators remain finite in the limit $c\to 0$; the rescaled generators vanish except for the surviving pair
@@
$g=\textnormal{d}x^{i}\otimes\textnormal{d}x^{j}\,\delta{ji}+2\,\textnormal{d}t\otimes\textnormal{d}\tau\,\tau-\textnormal{d}t\otimes\textnormal{d}t\,.$
@@
$\langle Q|\partial_{i}\rangle=\langle\partial\tau|\partial_{i}\rangle+\tau\,\langle\partial_{t}|\partial_{i}\rangle=0\,$,
@@
which demonstrates that the kernel of $g$ is generated by $Q$. Notice that the reduced metric is the Minkowski metric on ${\mathbb{R}}^{4}$.
Remark 1.1.
It must be stressed that the Inönü–Wigner contraction employed above does not yield a Carrollian structure as the underlying Poncaré group is not contracted. That is, we do not have a super-generalisation of a Carrollian structure, but rather an odd analogue thereof in that the fundamental vector field is now odd.
The above constructions will form the basis of examples 2.3 and 2.5.
The general definition of a super-null Riemannian manifold (see Definition 2.9) covers the above example; however, there will be no assumption that the supermanifold necessarily arises from an Inönü–Wigner contraction. That is, we take an intrinsic point of view.
@@
  • Main Theorem - affine connections that are both supersymmetric compatible and metric compatible exist on any super-null Riemannian manifold, and they carry non-vanishing torsion;
@@
  • Proposition 2.3 - the reduced metric is non-degenerate, that is, underlying a super-null Riemannian manifold is a pseudo-Riemannian manifold;
@@
  • Proposition 2.5 + Corollary 2.2 - the super-null Lie algebra is finite dimensional.
Arrangement. In Section 2 we proceed with the bulk of this paper. We recall the fundamental theory of supermanifolds equipped with (degenerate) metrics and connections in Subsection 2.1. In Subsection 2.2, the concept of a super-null Riemannian manifold is presented, and some immediate results are given. The super-null Lie algebra is studied in Subsection 2.3. The question of compatible affine connections is addressed in Subsection 2.4. We end in Section 3 with some concluding remarks.
@@ 2.1. Even Metrics and Connections on Supermanifolds
We will assume that the reader has some familiarity with the category of (real and finite-dimensional) supermanifolds SMan. We understand a supermanifold $M:=(|M|,:\mathcal{O}_{M})$ of dimension $n|m$ to be a supermanifold as defined by Berezin & Leites [3, 26], i.e., we take the locally ringed space approach. For an overview of the general theory of supermanifolds, the reader may consult, for example, [8, 25, 32]. Underlying any supermanifold is a smooth manifold that we will denote $M_{red}=(|M|,C^{\infty}_{|M|}(-))$. An incomplete list of works on Riemannian supermanifolds includes [12, 13, 14, 15, 17, 24]. The warning from the outset is that we will include degenerate metrics in our definition.
@@
for all (homogeneous) $X,Y,Z\in{Vect}(M)$ and $f\in C^{\infty}(M)$. Extension of these properties to inhomogeneous vector fields is by linearity. A vector field $X\in{Vect}(M)$ is said to be a null vector field if $\langle X|X\rangle=0$. Note that odd vector fields are automatically null vector fields as, via the symmetry of the inner product $\langle X|X\rangle=-\,\langle X|X\rangle=0$.
@@
A metric $g$ is said to be non-degenerate if $\ker(g)={\mathbf{0}}$, and is said to be degenerate otherwise.
@@
Killing vector fields are defined in the same way as in classical Riemannian geometry.
@@
This follows in complete parallel with the classical case using $\mathcal{L}_{[X,Y]}=[\mathcal{L}_{X},\mathcal{L}_{Y}]$. ∎
@@
Affine connections exist on any (real) supermanifold. This can be proved by adapting the standard arguments, i.e., affine connections are local operators and the existence of a partition of unity [^2] See [26, Lemma 3.1.7 and Corollary 3.1.8] for the existence of partitions of unity and bump functions on supermanifolds..
@@
2.2. Super-Null Riemannian Manifolds
@@
A super-null Riemannian manifold is a quadruple $(M,g,Q,P)$, where
@@
subject to the compatibility condition $\ker(g)=\operatorname{Span}{Q}$.
A morphism of super-null Riemannian manifolds $\Phi:(M,g,Q,P)\rightarrow(M^{\prime},g^{\prime},Q^{\prime},P^{\prime})$ is a diffeomorphism $\Phi:M\rightarrow M^{\prime}$, such that
@@
The resulting category of super-null Riemannian manifolds is denoted SNullMan. The group of automorphism of a super-null Riemannian manifold we denote as $\mathrm{SNull}(M,g,Q,P)$, and the associated super-null Lie algebra $\mathfrak{snull}(M,g,Q,P)$ is the Lie superalgebra of vector fields $X\in{Vect}(M)$ that satisfy
@@
    the rank of the kernel is constant, which ensures the kernel is well-defined as a locally free module;
@@
    the supermanifold $M$ we take to be of dimension $n|1$, with $n\geq 1$, this is minimal case and parallels $N=1$ SUSY-curves. Moreover, $M$ is an odd line bundle, i.e., locally $M$ is of the form $|U|\times{\mathbb{R}}^{0|1}$;
@@
    Shander’s theorem [30] means that in the neighbourhood of any point $m\in|M|$, adapted coordinates can always be found, which we will refer to as Shander coordinates, $(x^{a},t,\tau)$, such that the vector fields take the canonical form
@@
    morphisms equate the kernels, i.e., $\Phi^{*}\ker(g)=\ker(g^{\prime})$, and $\Phi^{*}P=P^{\prime}$ as the push-forward by a diffeomorphism is a homomorphism of Lie brackets;
@@
    if $X\in\mathfrak{snull}(M,g,Q,P)$, then $\mathcal{L}_{X}P=[[X,Q],Q]=0$.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold. Then in Shander coordinates $(x^{a},t,\tau)$, the most general form of the degenerate metric $g$ is
@@
In Shander coordinates $(x^{a},t,\tau)$, the most general form of a (possibly degenerate) metric on $M$ is of the form
@@
Locally, $\tau Q=\tau\partial_{\tau}$, and so these vector fields are not linearly independent. Thus, $\langle Q|\partial_{\tau}\rangle=0$ is consistent with the rank of $\ker(g)$ being generated by $Q$. Moreover, we have not yet placed constraints on the local components of the metric.
@@
which, a priori, may be a degenerate. Using coordinates on $M_{red}$ induced by Shander coordinates $(x^{a},t,\tau)$, the metric $g_{red}$ is of the form
@@
The reader can quickly check that we do indeed have a super-null Riemannian manifold, i.e., the kernel of $g$ is generated by $Q$. The reduced metric is $g_{red}=\textnormal{d}x\otimes\textnormal{d}x+\textnormal{d}t\otimes\textnormal{d}t$, which is non-degenerate.
@@
Let $(M,g,Q,P)$ be a super-null Riemannian manifold. Then the reduced manifold $M_{red}$ is a pseudo-Riemannian manifold.
@@
As the question of degeneracy can be addressed locally, we will employ Shander coordinates $(x^{a},t,\tau)$, so that $Q=\partial_{\tau}+\tau\partial_{t}$ and $P=\partial_{t}$. By Proposition 2.2, the reduced metric has the form
@@
$n=1$: Degenerate metrics are locally of the form
where $g_{tt}$ is a smooth function $t$. To have a constant rank kernel, it must be the case that $g_{tt}$ is nowhere vanishing. The reduced metric is $g_{red}=\textnormal{d}t\otimes\textnormal{d}t\,g_{tt}(t)$, which is non-degenerate.
$n\geq 2$: As a block matrix we have
$g_{red}=\begin{pmatrix}g_{ab}&g_{at}\\ g_{tb}&g_{tt}\end{pmatrix}\,$,
where the entries are ordinary smooth functions in $x^{a}$ and $t$. Whenever $g_{ab}$ is invertible, we may factor the determinant by the Schur complement:
$\det(g_{red})=\det(g_{ab})\cdot S\,,\qquad S:=g_{tt}-g_{ta},(g_{ab})^{-1}g_{bt}\,.$
@@
Claim 1: for $n\geq 2$, $\det(g_{ab})\neq 0$. We will prove this via a contradiction. Let us assume that there exists a $U=U^{a}\partial_{a}$ such that $U^{a}g_{ab}=0$. Then any vector field $X=X^{a}\partial_{a}+X^{t}\partial_{t}+X^{\tau}\partial_{\tau}$ we have
@@
Next, consider a new vector field $U^{\prime}=U+\lambda,P$, where $\lambda$ is an even function. Then
@@
Claim 2: for $n\geq 2$, The Schur scalar $S$ is nonzero. Claim 1 establishes that $g_{ab}$ is invertible, and we will denote the inverse as $g^{ab}$, as standard. The Schur scalar is $S=g_{tt}-g_{ta}g^{ab}g_{bt}$. If $S=0$, then there exists a vector field $V\in\ker(g_{red})$. Every such even vector field has a canonical lift, also denoted here by $V$, as it does not depend on the odd coordinate. Thus, $\epsilon(\langle V|X\rangle)=0$ for all lifts $X$, and so $\langle V|X\rangle=0$. Thus, $V$ is in the kernel of $g$. However, this violates the condition on the kernel’s rank. This implies it must be the case that $S\neq 0$.
Thus, for $n\geq 2$, $\det(g_{red})\neq 0$ and so $g_{red}$ defines a pseudo-Riemannian structure on $M_{red}$.∎
@@
where $g_{\mathbb{R}}$ is the constant metric on $\mathbb{R}$}. Then $M=M_{0}\times{\mathbb{R}}^{1|1}$ is a super-null Riemannian manifold with the degenerate metric in Shander coordinates being of the form
@@
2.3. The Super-Null Lie Algebra
The super-null Lie algebra was defined earlier as the infinitesimal automorphisms of a super-null Riemannian manifold (see Definition 2.9). The Lie superalgebra $\mathfrak{snull}(M,g,Q,P)$ is defined as the Lie superalgebra of vector fields $X\in{Vect}(M)$ that satisfy $\mathcal{L}_{X}g=0$ and $\mathcal{L}_{X}Q=0$. In this subsection, we examine the structure of this Lie superalgebra and establish that it is finite dimensional.